OpenStudy (anonymous):

solve the system of equations 2x+y=11

OpenStudy (pinksapphire):

Do you have the other equation?

OpenStudy (anonymous):

oh no hold up, that does not fit

OpenStudy (pinksapphire):

She has another equation with it...

OpenStudy (anonymous):

x-y=4 is the other equation with it.

OpenStudy (pinksapphire):

Solve for y first... 2x+y=11 -2x -2x y=11-2x Now plug in 11-2x in for y in the other equation... x-y=4 x-(11-2x)=4 Distribute... x-11+2x=4 +11 +11 x+2x=15 3x=15 Divide both sides by 3 3x/3=15/3 x=5 Plug in any equation... x-y=4 (5)-y=4 +5 +5 -y=9 Divide by -1 to get y to be a positive number... -y/-1=9/-1 y=-9 and x=5

OpenStudy (anonymous):

Substitution is the best method for this. Change x-y+4 to y=x-4 and plug in 2x+x-4=11 3x-4=11 3x=15 x=5 and if you plug that value in, 2(5)+y=11 y=1 so (5,1)

OpenStudy (anonymous):

would substitution method be the best also for this one to: 5x-2y=28 3x+47=22

OpenStudy (anonymous):

No, the elimination would be best. Do you want me to work it out for you?

OpenStudy (pinksapphire):

I think there is a mistake in the other equation that I did...

OpenStudy (anonymous):

hey did you mean 4y or 47?

OpenStudy (anonymous):

not 47 4y yes please

OpenStudy (anonymous):

okay. I will work it out as 4y

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since 4 is 2 times 2, multiply 2 with every thing on top equation :) 5x(2)-2y(2)=28(2) which would be 10x-4y=56 then, 10x-4y=56 (+) 3x+4y=22 ___________________ 13x=78 x=6 Plug in on either original equation. (I will go with bottom one) 3(6)+4y=22 18+4y=22 4y=4 y=1 So (6,1) check; 5(6)-2(1)=28 30-2=28 3(6)+4(1)=22 18+4=22 DONE!!!! :D

OpenStudy (anonymous):

thanky ou!!!!